OpenStudy (anonymous):

The speed of light is 3x10*8(* =to the power of), how far can light travel in one day in scientific notation ?

OpenStudy (anonymous):

speed of light needs units. \(3\times 10^8\) what? miles per hour? meters per second? you are going to have to convert the units to days, so we need to know it before beginning

OpenStudy (anonymous):

meters a second.

OpenStudy (anonymous):

One day = 24 * 60 * 60 seconds = 86400 seconds So, distance travelled in one day by light = 3 * 10^8 * 86400 = 25920000000000 = 2.592 * 10^13

OpenStudy (anonymous):

Distance = rate * time =D Awesome!

OpenStudy (anonymous):

ok so now what pratu said. you need to know how many seconds are in a day. 60 seconds in a minute, 60 minutes in an hour and 24 hours in a day. so multiply \(60\times 60\times 24\) and then multiply the result by \(3\times 10^8\)

OpenStudy (anonymous):

cheers guys